Write an equation of the line containing the specified point and parallel to the indicated line.
(-7,0),5x+2y=6

Answer by htmentor(1343)

The equation will be of the form y = mx + b
Putting the equation for the indicated line in this form gives
y = -5/2x + 3
The line we want is parallel to the above which means the slopes are the same.
So the equation of the line is y = -5/2x + b
The point (-7,0) lies on the line. We can use this information to solve for b.
0 = -5/2*-7 + b -> b = -35/2.
So the answer is:
y = -5/2x - 35/2
Here are the graphs of the two equations: